Complete the gaps in the following table:

Nominal APR

Number of Compounding Periods

Rate per Period

Equivalent Annual Rate

10.75%

52

0.207%

4

3.12%

12%

365

0.03287%
Answer #1

Case A:

Equivalent Annual Rate = (1 + Rate per Period)^Number of Compounding Periods - 1
Equivalent Annual Rate = (1 + 0.00207)^52 - 1
Equivalent Annual Rate = 1.1135 - 1
Equivalent Annual Rate = 0.1135 or 11.35%

Case B:

Nominal APR = Rate per Period * Number of Compounding Periods
Nominal APR = 3.12% * 4
Nominal APR = 12.48%

Equivalent Annual Rate = (1 + Rate per Period)^Number of Compounding Periods - 1
Equivalent Annual Rate = (1 + 0.0312)^4 - 1
Equivalent Annual Rate = 1.1308 - 1
Equivalent Annual Rate = 0.1308 or 13.08%

Case C:

Equivalent Annual Rate = (1 + Rate per Period)^Number of Compounding Periods - 1
Equivalent Annual Rate = (1 + 0.0003287)^365 - 1
Equivalent Annual Rate = 1.1274 - 1
Equivalent Annual Rate = 0.1274 or 12.74%